OpenStudy (anonymous):

8+2(8x-6)= Choose exactly two answers that are correct. A. 2(4x+7) B. 9x-10 C. 16x-4 D. 4(4x-1)

OpenStudy (anonymous):

C. and D. Expand D. and you will have C.

OpenStudy (anonymous):

can u help with more

OpenStudy (anonymous):

oh i can help you with more bb ;)

OpenStudy (jonnyvonny):

@robtobey Great helping, mate.

OpenStudy (anonymous):

Thank you @jonnyVonny

OpenStudy (jonnyvonny):

Obvious sarcasm.

OpenStudy (anonymous):

1.5x-20=2(3x-10) 2.4x-16-2x=2(x-8) 3.6x-15=3(2x-6) 4.5x-3(x-4)=2(x-6) What is the best way to classify each equation? A. identity B. contradiction C. neither

OpenStudy (anonymous):

thats right tootsie tits, keep asking more questions mmmm i love it when you get all photosynthesis with me

OpenStudy (anonymous):

WTF